# A Generative Marker Enhanced End-to-End Framework for Argument Mining

### v2

@@ Abstract
Index Terms:
Argument Mining, Marker, Generative Paradigm, Argument Component Classification, Argumentative Relation Classification, Augmented Natural Language.
@@ I Introduction
Argument Mining (AM) [1] deals with the detection and classification of ACs and their corresponding ARs from discourse dynamics. Figure 1 gives an illustrative example. AM task has been subdivided into four sub-tasks [2]: (i) Component Segmentation entails identifying Argumentative Discourse Units (ADUs) [3], (ii) Component Classification involves categorizing ADUs into various ACs [4], (iii) Relation Identification focuses on detecting argumentative relationships among two or more ADUs [5], and (iv) Relation Classification deals with classifying these identified relations into different ARs [6]. However, following the study in [7], we collectively term the initial pair of sub-tasks as Argument Component Extraction (ACE) and the subsequent pair as Argumentative Relation Classification (ARC). Consequently, end-to-end AM, referred to as ACRE in this work, involves jointly addressing both ACE and ARC tasks. AM is useful for various downstream tasks such as debate analysis [8], automated essay scoring [9], customer review analysis [10], etc.
Figure 1: An illustrative example of Argumentative Structure with Argumentative Components (ACs: Claim and Premise) and Argumentative Relations (ARs: Support and Attack). Premise 1 supports the claim, whereas, Premise 2 attacks the claim.
@@
The primary challenge in AM lies in effectively handling the longer sequence length of ACs and their associated ARs [1]. Defining boundaries for ACs is more intricate compared to tasks like Named Entity Recognition (NER) or Parts-of-Speech (POS) tagging, where the target text span consists of a few tokens only [11]. Also, every AC has certain underlying contexts of argumentativeness and is related to another AC of the same context. Variations in argument representations across domains pose another challenge [12]. Previous approaches to this task have primarily utilized extractive methods [13] to address these challenges, which have shown limited performance. Recently, end-to-end frameworks based on dependency parsing [7] have gained popularity due to their ability to streamline the extraction process into a single unified model. However, these methods often require intricate post-processing steps, adding to the overall complexity of the solution. Given the limitations of these recent approaches, we propose an alternative method based on a text-to-text generation paradigm. This new approach reduces complexity by eliminating the need for elaborate post-processing.
@@
Several generative frameworks have recently been proposed for various NLP tasks, including NER [14], joint entity and relation extraction [15], and co-reference resolution [16]. One such framework is Translation between Augmented Natural Language (TANL)[17], which addresses a range of NLP tasks within a unified framework. This framework generates outputs by augmenting the original input text with class labels specific to downstream tasks. “[ Tolkien |||| person ]’s epic novel [ The Lord of the Rings |||| book ] was published in 1954-1955.” Here, labels such as person and book are embedded within the original text. With this kind of output format, complex post-processing steps are reduced. However, among all the NLP tasks solved by the TANL framework, the target output spans (e.g., Tolkien or The Lord of the Rings) are relatively shorter as compared to the target spans of AM tasks, where each AC spans across longer sections of text. Thus, it is a meaningful avenue to explore the capability of TANL in solving AM tasks that consist of longer target spans.
@@
Studies [18, 19] have demonstrated that markers are strongly correlated with improvements in AM task performance as they reliably indicate the presence of ACs in the text. Two distinct types of markers are categorized: (a) Argumentative Markers and (b) Discourse Markers (DMs) (See Figure 3). Argumentative Markers are typically phrases such as “I strongly agree that,” “But, I deny the point that,” or “However, this clearly proves that,” conveying the argumentativeness of the discourse. In contrast, DMs are single-token connectives like “But,” “And,” or “However,” representing the rhetorical structure of language. The set of argumentative markers is extracted using the rule-based methods by [20], while the set of DMs is constructed from a specific web-corpus by [21]. However, rule-based extraction yields a large number of irrelevant spans lacking generalizability. We employ a similar extraction method for Argumentative Markers, incorporating an additional manual filtering step to eliminate irrelevant or non-generic spans.
@@
We conduct a comprehensive set of experiments to evaluate the performance of the proposed framework, organized into several subsections. We first analyze the main results with a detailed comparison of argTANL, Marker-Based Fine-Tuned argTANL, and ME-argTANL to highlight the improvements brought by the knowledge of markers. Then we compare the Component-Only and End-to-End variants to determine their relative effectiveness. Next, we assess the results of the Relation-Only variant to understand its standalone performance. We explore Coarse-Grained vs. Fine-Grained Label Semantics to determine how label-granularity impacts the performance of the proposed framework. We also analyze the Impact of input length on the performance and investigated the framework’s ability to handle Relations of different depths. To evaluate the argTANL framework, we conduct an ablation study by introducing Abbreviated argTANL. Here, repeated full AC spans are replaced with unique “ID” tokens, reducing output sequence length and improving efficiency in handling longer paragraphs. This study explores the performance trade-offs between the standard and abbreviated formulations. Finally, an Error Analysis is performed to find the scope for further improvement. Through these extensive experiments upon three standard benchmarks of AM literature, the proposed approach achieves State-of-the-Art (SoTA) results in several datasets and surpasses various important baselines. In summary, the main contributions of this paper are:
Figure 3: Illustrative instances of Argumentative and Discourse markers present in the standard AM corpus.
@@
    A generative task formulation in the form of argTANL framework to solve End-to-End AM along with Component-only and Relation-only variants.
@@
II Task Formulation
@@
We represent argumentative text as $W=w_{1},w_{2},w_{3},....,w_{n}$, where $n$ is the total number of tokens in $W$. For a text-span $w_{i},w_{i+1},w_{i+2},\ldots,w_{j}$ in $W$, we write it as $w_{i:j}$. We define a set of AC types as $T^{c}=\{t_{1}^{c},t_{2}^{c},t_{3}^{c},....,t_{n_{c}}^{c}\}$ and a set of AR types as $T^{r}=\{t_{1}^{r},t_{2}^{r},t_{3}^{r},....,t_{n_{r}}^{r}\}$, where $n_{c}$ and $n_{r}$ refer to a total number of possible AC and AR types respectively. We use a set of symbol tokens, $S=\{[,],=,|\}$ in the ANL, where ``["``["` ` [ " marks the start of a component, ``]"``]"` ` ] " marks the end of a component, ``="``"``="` ` = " is used for relation assignments, and ``|"conditional``"``|"` ` | " is used to separate labels. Subsequent sections discuss different formulations of the proposed task.
II-A ACE task: Component-only Variant
@@
For any given argumentative text $w_{1:n}$, objective of the ACE is to extract a set of ACs as $C=\{C_{i}|C_{i}=(c_{i},c_{i}^{s},c_{i}^{e})\}$, where $C_{i}$ is the $i^{th}$ AC, $c_{i}\in T^{c}$, and $c_{i}^{s}$ and $c_{i}^{e}$ refer to the relative start and end token indices of $c_{i}$ respective to $W$. Here, we generate the label-augmented text for ACs only such as, in Figure 2, for the head, AC is [ the endangered species preserved in zoos would never die of illegal hunting |||| Premise ] and for tail AC is [ zoos, which are equipped with modern facilities and professionals, would provide better care for the animals inside |||| Claim ]. The main motivation for this variant is to assess our proposed method’s ability to identify ACs alone when the relational information is muted.
II-B ARC task: Relation-Only Variant
It differs in terms of the given input text and the target output text. Here, the input is also an ANL with an indication of ACs’s boundaries without the corresponding type information. The target output is also an ANL of a specified format: “Relationship between [ADU 1] and [ADU 2] is = Relation-Type”. Notably, at any point in time, only two ADUs are taken to form an output. For example, in Figure 2, input ANL is “Nevertheless, supporters would argue that [ zoos, which are equipped … care for the animals inside ]. They further point out that [ the endangered species … illegal hunting ].”. The corresponding output ANL is “Relationship between [zoos, which are equipped … care for the animals inside ] and [ the endangered species … illegal hunting ] is = Support”.
II-C ACRE task: End-To-End Argument Mining
@@
The proposed end-to-end formulation jointly frames the ACE and ARC tasks in the following manner. We define ARs as $R=\{R_{i}|R_{i}=(r_{i},r_{i}^{h},r_{i}^{t})\}$, where $R_{i}$ is the $i^{th}$ AR corresponding to AR type $r_{i}\in T^{r}$, and $r_{i}^{h}$ and $r_{i}^{t}$ refer to the head and tail ACs respectively. $r_{i}^{h}$ and $r_{i}^{t}$ are connected with relation type $r_{i}$. If two components $C_{p},C_{q}\in C$ are related with $R_{k}\in R$ as $r_{k}^{h}=C_{p}$ and $r_{k}^{t}=C_{q}$, then for the token spans of $C_{p}$ i.e. $w_{c_{p}^{s}:c_{p}^{e}}$ and $C_{q}$ i.e. $w_{c_{q}^{s}:c_{q}^{e}}$, the model will generate augmented labels as $[w_{c_{p}^{s}:c_{p}^{e}}|c_{p}|r_{k}=w_{c_{q}^{s}:c_{q}^{e}}]$ and $[w_{c_{q}^{s}:c_{q}^{e}}|c_{q}]$ respectively. The rest of the tokens in $W$ will be rewritten as it is. We refer to this joint formulation as ACRE. Figure 2 illustrates the ACRE formulation.
III Methodology
@@
III-A The argTANL Framework
This framework adopts the TANL formulation (Figure 2) with specific customization tailored for the AM task. Given argumentative text as input, the framework generates ANL consisting of ACs and ARs as label-augmented texts enclosed with symbol tokens, as described in Section II. This approach leverages the strengths of the TANL framework while introducing modifications to effectively capture the complex relationships and structures inherent in argumentative texts. The integration of symbolic tokens ensures precise boundary definition and clear distinction between different ACs and ARs, facilitating accurate identification within the generated sequences.
TABLE I: Examples of extracted valid markers and manually filtered-out topic-dependent marker candidates (non-markers) from AAE corpus, which are common in our extracted list of marker candidates and the marker list provided by [20]. Topic information is marked in bold.
Valid Markers
(i) Nevertheless, I believe that
(ii) Another supporting reason is that
(iii) People who hold different opinion may argue that
(iv) I strongly disagree with this affirmation because I believe
(v) In conclusion, the above stated reasons clearly outweigh the fact that
Manually Filtered-Out Topic-Dependent Marker Candidates
(i) In spite of the importance of sports activities
(ii) Moreover, the proponents of globalization idea point out
(iii) Nevertheless, opponents of online-degrees would argue that
(iv) The official term of it is named “technological unemployment”
(v) However, as the society grows, human rights become more highly respected
III-B Argumentative Marker Extraction
An argumentative marker typically signals the beginning of an AC. However, not every marker is always followed by an AC, and an AC may not always be preceded by a marker. Considering this phenomenon, we extract two types of (See Figure 4) potential marker candidates from any argumentative text: (i) Leading candidates: by extracting tokens from the start of a sentence to the beginning of an AC, and (ii) Sandwich candidates: by extracting tokens after the end of an AC until the start of another AC in the same sentence. However, these strategies yielded some spans, which were non-generic. For example, in Table I, “In spite of the importance of sports activities” or “Nevertheless, opponents of online-degrees would argue that” are having non-generic spans. These spans lacked generalizability across diverse topics, being tailored solely to their respective subjects. To enhance generalizability, we introduce an additional layer of manual filtering to eliminate topic-dependent spans from the pool of extracted marker candidates. Initially, we identified 2925 marker candidates. After manual filtering and removing duplicates, we refine the list to 1072 unique argumentative markers.
TABLE II: Description of the initial step of Marker-Based Fine-Tuning of the argTANL framework.
@@ III-B Argumentative Marker Extraction
A-MKT Last but not least, students have … difficulties. [ Last but not least, |||| marker ] students have … difficulties.
SM-MKT $<$extra_id_0$>$ $<$extra_id_1$>$ $<$extra_id_2$>$ $<$extra_id_3$>$ $<$extra_id_4$>$ students have … difficulties. $<$extra_id_0$>$ Last $<$extra_id_1$>$ but $<$extra_id_2$>$ not $<$extra_id_3$>$ least $<$extra_id_4$>$ , $<$extra_id_5$>$
@@
III-C Marker-Based Fine-Tuning of argTANL
This technique strategically divides the argTANL output generation into two significant steps. The initial step leverages the extracted markers from various corpora to execute Marker-Based Fine-Tuning. This involves implementing four distinct generative fine-tuning strategies, each utilizing varied input and output combinations (see Table II). The objective is to acquaint the model with the nuanced representations of markers within the argumentative text. Notably, neither the markers from the test data nor the test data itself were used during this initial fine-tuning process.
@@ III-C Marker-Based Fine-Tuning of argTANL
III-C1 Augmented Marker Knowledge Transfer (A-MKT)
@@
III-C2 Span-Masked Marker Knowledge Transfer (SM-MKT)
@@
III-C3 Marker Knowledge Transfer through Encoding (E-MKT)

### v3

# Nilmadhab Das, Vishal Choudhary, V. Vijaya Saradhi, Ashish Anand

@@ I Introduction
Argument Mining (AM) [1] deals with the detection and classification of ACs and their corresponding ARs from discourse dynamics. AM task has been subdivided into four sub-tasks [2]: (i) Component Segmentation entails identifying Argumentative Discourse Units (ADUs) [3], (ii) Component Classification involves categorizing ADUs into various ACs [4], (iii) Relation Identification focuses on detecting argumentative relationships among two or more ADUs [5], and (iv) Relation Classification deals with classifying these identified relations into different ARs [6]. However, following the study in [7], we collectively term the initial pair of sub-tasks as Argument Component Extraction (ACE) and the subsequent pair as Argumentative Relation Classification (ARC). Consequently, end-to-end AM, referred to as ACRE in this work, involves jointly addressing both ACE and ARC tasks. AM is useful for various downstream tasks such as debate analysis [8], automated essay scoring [9], customer review analysis [10], etc.
@@
The primary challenge in AM lies in effectively handling the longer sequence length of ACs and their associated ARs [1]. Defining boundaries for ACs is more intricate compared to tasks like Named Entity Recognition (NER) or Parts-of-Speech (POS) tagging, where the target text span consists of a few tokens only. Also, every AC has certain underlying contexts of argumentativeness and is related to another AC of the same context. Variations in argument representations across domains pose another challenge [11]. Previous approaches to this task have primarily utilized extractive methods [12] to address these challenges, which have shown limited performance. Recently, end-to-end frameworks based on dependency parsing [7] have gained popularity due to their ability to streamline the extraction process into a single unified model. However, these methods often require intricate post-processing steps, adding to the overall complexity of the solution. Given the limitations of these recent approaches, we propose an alternative method based on a text-to-text generation paradigm. This new approach reduces complexity by eliminating the need for elaborate post-processing.
@@
Several generative frameworks have recently been proposed for various NLP tasks, including NER [13], joint entity and relation extraction [14], and co-reference resolution [15]. One such framework is Translation between Augmented Natural Language (TANL)[16], which addresses a range of NLP tasks within a unified framework. This framework generates outputs by augmenting the original input text with class labels specific to downstream tasks. “[ Tolkien $|$ person ]’s epic novel [ The Lord of the Rings $|$ book ] was published in 1954-1955.” Here, labels such as person and book are embedded within the original text. With this kind of output format, complex post-processing steps are reduced. However, among all the NLP tasks solved by the TANL framework, the target output spans (e.g., Tolkien or The Lord of the Rings) are relatively shorter as compared to the target spans of AM tasks, where each AC spans across longer sections of text. Thus, it is a meaningful avenue to explore the capability of TANL in solving AM tasks that consist of longer target spans.
@@
Studies in [17] have demonstrated that markers are strongly correlated with improvements in AM task performance as they reliably indicate the presence of ACs in the text. Two distinct types of markers are categorized: (a) Argumentative Markers and (b) Discourse Markers (DMs). Argumentative Markers are typically phrases such as “I strongly agree that,” “But, I deny the point that,” or “However, this clearly proves that,” conveying the argumentativeness of the discourse. In contrast, DMs are single-token connectives like “But,” “And,” or “However,” representing the rhetorical structure of language. The set of argumentative markers is extracted using the rule-based methods by [18], while the set of DMs is constructed from a specific web-corpus by [19]. However, rule-based extraction yields a large number of irrelevant spans lacking generalizability. We employ a similar extraction method for Argumentative Markers, incorporating an additional manual filtering step to eliminate irrelevant or non-generic spans.
@@
We conduct a comprehensive set of experiments to evaluate the performance of the proposed framework, organized into several subsections. We first analyze the main results with a detailed comparison of argTANL, Marker-Based Fine-Tuned argTANL, and ME-argTANL to highlight the improvements brought by the knowledge of markers. Then we compare the Component-Only and End-to-End variants to determine their relative effectiveness. Next, we analyze the Impact of input length on the performance of the argTANL. Then, we conduct an ablation study by introducing Abbreviated argTANL. Here, repeated full AC spans are replaced with unique “ID” tokens, reducing output sequence length and improving efficiency in handling longer paragraphs. This study explores the performance trade-offs between the standard and abbreviated formulations. Finally, an Error Analysis is performed to find the scope for further improvement. Through these extensive experiments upon three standard benchmarks of AM literature, the proposed approach achieves competitive results in several datasets and surpasses various important baselines. In summary, the main contributions of this paper are:
@@
    A generative task formulation in the form of argTANL framework to solve End-to-End AM along with Component-only variant.
@@
II Related Work
Argument Mining: Prior studies largely focus on subsets of the four AM sub-tasks. Recent works [20, 12] emphasize joint formulations in an end-to-end manner. [21] adopted a pipelined approach for ACE and ARC, optimizing error propagation via Integer Linear Programming (ILP). [22] reformulated the end-to-end AM task into sequence tagging, dependency parsing, multi-task tagging, and relation extraction. [20] proposed a biaffine dependency parsing approach for AM. Dataset scarcity was highlighted by [7], who introduced a cross-corpora multi-task span-biaffine framework, employing a span classifier with BIO tagging and span representations via average pooling. [12] framed AM as a text-to-sequence generation task, producing sequences of AC and AR types with start/end indices. Recently, the TANL framework was used by [23] to address end-to-end AM as text-to-text generation, but exploration of markers and ANL formats remained limited. Our work addresses this gap.
Markers: Markers are pivotal signals for ADUs [24]. [25] used markers as lexical features for classifying argument components via multiclass classification. [18] extracted 1131 argumentative markers to improve span representations in AM tasks. [24] explored markers in Reddit threads, extracting 69 Reddit-specific markers and applying selective masked language modeling (sMLM) for domain adaptation. The model was then used for Argument Component Identification, with marker-like tokens predicted to infer Relation Types. However, marker exploration for AM in a generative, end-to-end framework remains unexplored.
Augmented Natural Language (ANL) & Generative Paradigm: Generative methods are reframing NLP tasks as generation problems, with label-augmented text (ANL) being a notable strategy. ANL has been applied to NER [26], sentiment analysis [27], and relation extraction [28]. [16] employed ANL for structured prediction tasks such as joint entity and relation extraction, event argument extraction, and coreference resolution by framing them as text-to-text generation problems.
III Task Formulation
@@
We represent argumentative text as $W=w_{1},w_{2},w_{3},....,w_{n}$, where $n$ is the total number of tokens in $W$. For a text-span $w_{i},w_{i+1},w_{i+2},\ldots,w_{j}$ in $W$, we write it as $w_{i:j}$. We define a set of AC types as $T^{c}=\{t_{1}^{c},t_{2}^{c},t_{3}^{c},....,t_{n_{c}}^{c}\}$ and a set of AR types as $T^{r}=\{t_{1}^{r},t_{2}^{r},t_{3}^{r},....,t_{n_{r}}^{r}\}$, where $n_{c}$ and $n_{r}$ refer to a total number of possible AC and AR types respectively. We use a set of symbol tokens, $S=\{[,],=,|\}$ in the ANL, where [``[" marks the start of a component, ]``]" marks the end of a component, =``=" is used for relation assignments, and |``|" is used to separate labels. Subsequent sections discuss different formulations of the proposed task.
III-A ACE task: Component-only Variant
@@
For any given argumentative text $w_{1:n}$, objective of the ACE is to extract a set of ACs as $C=\{C_{i}|C_{i}=(c_{i},c_{i}^{s},c_{i}^{e})\}$, where $C_{i}$ is the $i^{th}$ AC, $c_{i}\in T^{c}$, and $c_{i}^{s}$ and $c_{i}^{e}$ refer to the relative start and end token indices of $c_{i}$ respective to $W$. Here, we generate the label-augmented text for ACs only such as, in Figure 1, for the head, AC is [ the endangered species preserved in zoos would never die of illegal hunting $|$ Premise ] and for tail AC is [ zoos, which are equipped with modern facilities and professionals, would provide better care for the animals inside $|$ Claim ]. The main motivation for this variant is to assess our proposed method’s ability to identify ACs alone when the relational information is muted.
III-B ACRE task: End-To-End Argument Mining
@@
The proposed end-to-end formulation jointly frames the ACE and ARC tasks in the following manner. We define ARs as $R=\{R_{i}|R_{i}=(r_{i},r_{i}^{h},r_{i}^{t})\}$, where $R_{i}$ is the $i^{th}$ AR corresponding to AR type $r_{i}\in T^{r}$, and $r_{i}^{h}$ and $r_{i}^{t}$ refer to the head and tail ACs respectively. $r_{i}^{h}$ and $r_{i}^{t}$ are connected with relation type $r_{i}$. If two components $C_{p},C_{q}\in C$ are related with $R_{k}\in R$ as $r_{k}^{h}=C_{p}$ and $r_{k}^{t}=C_{q}$, then for the token spans of $C_{p}$ i.e. $w_{c_{p}^{s}:c_{p}^{e}}$ and $C_{q}$ i.e. $w_{c_{q}^{s}:c_{q}^{e}}$, the model will generate augmented labels as $w_{c_{p}^{s}:c_{p}^{e}}|c_{p}|r_{k}=w_{c_{q}^{s}:c_{q}^{e}}]$ and $w_{c_{q}^{s}:c_{q}^{e}}|c_{q}]$ respectively. The rest of the tokens in $W$ will be rewritten as it is. We refer to this joint formulation as ACRE. Figure 1 illustrates the ACRE formulation.
IV Methodology
@@
IV-A The argTANL Framework
This framework adopts the TANL formulation (Figure 1) with specific customization tailored for the AM task. Given argumentative text as input, the framework generates ANL consisting of ACs and ARs as label-augmented texts enclosed with symbol tokens, as described in Section III. This approach leverages the strengths of the TANL framework while introducing modifications to effectively capture the complex relationships and structures inherent in argumentative texts. The integration of symbolic tokens ensures precise boundary definition and clear distinction between different ACs and ARs, facilitating accurate identification within the generated sequences.
TABLE I: Description of the initial step of Marker-Based Fine-Tuning of the argTANL framework.
@@ IV-A The argTANL Framework
A-MKT Last but not least, students have … difficulties. [ Last but not least, $|$ marker ] students have … difficulties.
SM-MKT <<extra_id_0>> <<extra_id_1>> <<extra_id_2>> <<extra_id_3>> <<extra_id_4>> students have … difficulties. <<extra_id_0>> Last <<extra_id_1>> but <<extra_id_2>> not <<extra_id_3>> least <<extra_id_4>> , <<extra_id_5>>
@@
IV-B Argumentative Marker Extraction
An argumentative marker typically signals the beginning of an AC. However, not every marker is always followed by an AC, and an AC may not always be preceded by a marker. Considering this phenomenon, we extract two types of (See Figure 2) potential marker candidates from any argumentative text: (i) Leading candidates: by extracting tokens from the start of a sentence to the beginning of an AC, and (ii) Sandwich candidates: by extracting tokens after the end of an AC until the start of another AC in the same sentence. However, these strategies yielded some spans, which were non-generic. For example, “In spite of the importance of sports activities” or “Nevertheless, opponents of online-degrees would argue that” are having non-generic spans. These spans lacked generalizability across diverse topics, being tailored solely to their respective subjects. To enhance generalizability, we introduce an additional layer of manual filtering to eliminate topic-dependent spans from the pool of extracted marker candidates. Initially, we identified 2925 marker candidates. After manual filtering and removing duplicates, we refine the list to 1072 unique argumentative markers.
IV-C Marker-Based Fine-Tuning of argTANL
This technique strategically divides the argTANL output generation into two significant steps. The initial step leverages the extracted markers from various corpora to execute Marker-Based Fine-Tuning. This involves implementing four distinct generative fine-tuning strategies, each utilizing varied input and output combinations (see Table I). The objective is to acquaint the model with the nuanced representations of markers within the argumentative text. Notably, neither the markers from the test data nor the test data itself were used during this initial fine-tuning process.
@@ IV-C Marker-Based Fine-Tuning of argTANL
IV-C1 Augmented Marker Knowledge Transfer (A-MKT)
@@
IV-C2 Span-Masked Marker Knowledge Transfer (SM-MKT)
@@
IV-C3 Marker Knowledge Transfer through Encoding (E-MKT)